# A Survey on Adversarial Contention Resolution

### v3

@@ 3.1. AQT and Contention Resolution
Bienkowski et al. (bienkowski:distributed, ; BienkowskiJKK12, ) consider a model where each participant has a queue, and messages may include control bits. The adversary considered here differs from the leaky-bucket and window adversaries discussed earlier: at any time, an arbitrary number of packets can be injected into the queue of any participant. The metrics of interest are the total packets in all queues, called total load, and the maximum queue size. The authors prove that their deterministic algorithm achieves a total load that is an additive $O(n^{2})$ amount larger than the best offline algorithm, and the maximum queue size is at most an $n$-factor larger.
@@
Summary. There are several significant messages conveyed by the literature surveyed above. Randomized backoff algorithms continue to disappoint in that they are stable only for sub-constant injection rates. Interestingly, in terms of latency, deterministic algorithms can achieve performance that rivals and even surpasses traditional randomized backoff algorithms. However, most of these results make assumptions about unique IDs and/or a priori knowledge of $n$. The exception is the algorithm by Anantharamu and Clebus (anantharamu2015broadcasting, ), which assumes a $1$-activating adversary. This leads to a more general message: while stability and fairness seem like natural qualities that we desire in algorithm design, it is often not possible to achieve both simultaneously when there is even a little (i.e., constant) burstiness. Therefore, all of this paints a fairly dismal picture of what is possible under window and leaky bucket adversaries, especially for deterministic algorithms, given the many impossibility results that apply.
What avenues are there for positive results? As we will discuss later in Section 3.3, new randomized approaches can offer better performance. However, fundamentally, perhaps the AQT model is overly pessimistic; this has been investigated in more general settings (e.g., (BKS14, ).) Here, focusing on whether queue sizes can become unbounded under a very specific injection pattern, and under an infinite stream of packets, may lead to results that are artificially bleak. In the next section, we turn our attention to work that is concerned with a finite stream of packets that arrive under an arbitrary schedule.
@@ 3.2. Deterministic Approaches
In this section, we depart from parameterized adversaries and notions of stability, while still focusing on deterministic contention resolution algorithms. Here, our performance metric is latency (recall Section 1.1.3). Much of the contention-resolution work in this setting is referred to as $k$-selection, which is described as follows. There are $n$ participants of which $k$, for $1\leq k\leq n$, activate over time, each with a single packet to transmit. The activation times are selected by an adversary. As before, most of the literature operates in a model where each participant has a unique ID. Furthermore, $n$ is typically assumed to be known; throughout this section, we assume this to be true unless explicitly stated otherwise. The goal is to minimize the latency (recall from Section 1.1 ) required for all $k$ participants to successfully send their respective packets.
@@
Many results put a spin on John von Neumann’s quote by generating deterministic results from randomized approaches. Specifically, a common technique in this area is to design a randomized algorithm, and then use the probabilistic method (alon2016probabilistic, ) to argue that a deterministic analog exists. Here, we focus on the result by De Marco and Kowalski (doi:10.1137/140982763, ); however, related approaches are used in several other papers to give existence proofs for efficient transmission schedules (DBLP:conf/icdcs/MarcoKS19, ; de2013contention, ; DEMARCO20171, ; Chlebus:2016:SWM:2882263.2882514, ; ChlebusK04, ; ChlebusGKR05, ).
De Marco and Kowalski (doi:10.1137/140982763, ) consider the dynamic $k$-selection problem where there is a global clock, $k$ is unknown, and no-CD. Their main result is a non-adaptive, deterministic algorithm for contention resolution with latency $O(k\log(n)\log\log(n))$. Notably, this result almost matches a previous lower bound by Greenberg and Winograd (greenberg1985lower, ) in $\Omega(k\log(n)/\log k)$ for adaptive algorithms under the static case with collision detection.
@@
Finally, note that by the sum of a geometric series, the number of slots in $\mathcal{M}$ over which any participant executes prior to reaching row $\lg(k)+1$ is $O(k\log n\log\log n)$. Adding to this the (asymptotically smaller) $O(k\log(n\log n))$ slots for the Komlós and Greenberg result yields a total latency of $O(k\log n\log\log n)$.
@@ 3.2.1. Deterministic $k$-Selection
Greenberg and Winograd (greenberg1985lower, ) consider a multiple access channel with CD, for the two cases where $k$ is known and unknown. The authors prove a lower bound for deterministic algorithms of $\Omega(k\log n/\log k)$ in the static case, which of course applies to the dynamic case. This is contrasted against the upper bounds, claiming that the deterministic algorithms in Capetanakis (capetanakis:generalized, ), Hayes (hayes:adaptive, ), and Tsybakov and Mikhailov (tsybakov1978free-russian, ; tsybakov1978free, ) provide a nearly-tight latency of $\Theta(k+k\log(n/k))$.
With regard to upper bounds, De Marco and Kowalski (doi:10.1137/140982763, ) consider the case with a global clock and no-CD, as discussed above. Subsequently, De Marco et al. (DBLP:conf/icdcs/MarcoKS19, ) examine the power of acknowledgement-based versus non-acknowledgement-based deterministic non-adaptive approaches when there is no global clock and under the no-CD setting. For unknown $k$, they give a non-acknowledgement-based algorithm with latency $O(k^{2}\log n)$, and show that this can be improved to $O(k^{2}\log n/\log k)$ with an acknowledgment-based algorithm. These results are (nearly) tight; the authors also provide lower bounds on latency of $\Omega(k^{2}/\log k)$ for both cases. This illustrates a startling contrast with the case for known $k$, where the authors give an acknowledgement-based algorithm with $O(k\log k\log n)$ latency, which is known to be nearly tight (ClementiMS03, ).
@@ 3.2.2. Deterministic Wakeup
Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider a multi-channel scenario with no-CD and no global clock. Here, all stations may access all $b\geq 1$ channels at once, and with the same parameter $k\geq 1$ being an upper bound on the number of stations that become active over time. The authors give a non-adaptive, deterministic algorithm for the cases where $k$ is unknown that has latency $O(k\log^{1/b}k\log n)$. When $b=\Omega(\log\log n)$, a faster deterministic algorithm is possible, with delay $O((k/b)\log n\log(b\log n))$, and a lower bound of $\Omega((k/b)\log(n/k))$ is shown for deterministic algorithms. For ease of presentation, we note that the authors also consider jamming (a topic discussed further in Section 4), where each channel is jammed independently and uniformly with known probability $p$ per round, where $10<p<$ 1. Here, Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) give two deterministic algorithms that succeed with probability $1-O(1/\texttt{poly}(n))$, the first with latency $O(k\log(n)\log^{1/b}k/\log(1/p))$, and the second with latency $O((k/b)\log(n)\log(b\log(n))/\log(1/p))$ when $b=\Omega(\log(b\log n))$.
@@
Finally, under the SINR channel model, Jurdziński et al. (jurdzinski2013distributed, ) give a deterministic algorithm with $O(\log^{2}n)$ makespan when there is a global clock; however, without a global clock, the authors show a $\Omega(n)$ makespan. Complementing this result, Jurdziński and Stachowiak (jurdzinski2015cost, ) show that any algorithm with access to a global clock must incur $\Omega(\log^{2}n)$ makespan in the classical (non-SINR) channel model.
Summary. Much of the theoretical territory has been mapped out with some nearly-matching upper and lower asymptotic bounds on latency, and the results are less bleak than those surveyed under the AQT model. Notably, with no-CD and assuming a global clock, we see this for both $k$-selection (De Marco and Kowalski (doi:10.1137/140982763, )) and for wakeup (De Marco and Kowalski (DEMARCO20171, ; de2013contention, )), where the latter result depends on some limited system information.
Despite providing a rosier picture, deterministic approaches for contention resolution fall short of being able to offer asymptotically-optimal latency in the dynamic case; this is unsurprising given the analogous situation in the static case (Section 2.2). Similarly, deterministic solutions to the wakeup problem yield high latency in cases without a global clock, or where key parameters are unknown. A natural place to look for improvement—perhaps even asymptotic optimality—is the domain of randomized algorithms, which we survey next.
@@ 3.3. Randomized Approaches for Contention Resolution
When a packet $u$ arrives, it listens to the channel to determine whether it is already in use by other packets. Any packets that have arrived ahead of $u$ may still be attempting to succeed and, if so, they should also be emitting a “busy signal”, which we describe momentarily. If $u$ detects this busy signal, then it listens until this signal ceases, at which point it can start attempting to transmit. In this way, there are always two groups of packets: (1) active packets that are trying to transmit, while maintaining a busy signal, and (2) idle packets that are waiting for their turn on the channel. Given this isolation, active packets can execute a contention-resolution algorithm for the static case in order to succeed; afterwards, any idle packets can become active and repeat this process.
@@
The busy signal can take different forms. For example, given CD and a global clock, active packets can emit a ”busy” signal in each even-numbered slot; see Figure 6. Any idle packet that hears this signal or a collision infers that active packets are present and keeps waiting. However, having all packets send in every second slot may not be desirable, since sending consumes energy (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ). To reduce this cost, packets may emit the busy signal probabilistically, which can greatly reduce the sending cost; however, there is now a probability of error in the situation where some packets may still be active, but fail to emit a busy signal. When this happens, any idle packets will become active prematurely and interfere with the already-active packets.
@@
This interference is mitigated differently, depending on the setting. Bender et al. (BenderKPY18, ; bender:contention, ) assume CD and $O(\log n)$ control bits (recall Section 1.1.4) in each transmission on the shared channel. The busy signal occurs probabilistically every $O(1)$ slots, with each active packet sending an expected $O(1)$ times to maintain the signal. Interleaved with the busy signal are slots used to estimate the number of active packets, $m\leq n$, which allows them to execute TSTB (recall Section 2.1) starting with a window of size $\Theta(m)$. Additionally, a leader is elected from the set of active packets to broadcast how many slots remain in the execution of TSTB; this serves as a busy signal during the execution of TSTB. The leader is frequently changed so that each packet is responsible for $O(1)$ transmissions in expectation. Newly-arrived idle packets need only listen for a handful of slots to receive this information, which allows them to (i) avoid becoming active prematurely and (ii) avoid any further listening until the current execution of TSTB is finished. While the above process can fail—for example, the estimate is poor, the busy signal fails, or active packets do not succeed in the execution of TSTB—this occurs with probability at most polynomially small in the number of active packets, which allows for $O(n)$ expected latency. Additionally, the expected number of channel accesses performed per packet is $O(\log\log^{*}n)$. This low number of accesses translates is important to modern networks where devices are battery-powered, given that listening and sending on the channel consumes limited energy.
@@ 3.3.1. Randomized $k$-Selection
De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ) explore randomized solutions with no-CD. When $n$ is known (or a known to within a constant multiplicative factor), they provide a non-adaptive algorithm with $O(n)$ latency, which is asymptotically optimal. For the case where $n$ is unknown, a lower bound of $\Omega(n\log n/(\log\log n)^{2})$ latency is shown to hold for any non-adaptive, randomized algorithm; thus, there is an asymptotic separation in latency between adaptive and non-adaptive randomized algorithms. Additionally, for $n$ unknown, the authors provide a strongly non-adaptive, non-acknowledgment-based algorithm with $O(n\log^{2}n)$ latency; this latency can be improved to $O(n\log n/(\log\log n)^{2})$ with an acknowledgment-based algorithm, which comes close to matching their lower bound. Lastly, when $n$ is unknown, the authors provide an adaptive algorithm that achieves $O(n)$ latency. Informally, this algorithm employs LE to break the dynamic problem into a sequence of static contention-resolution problems, as discussed earlier in this section.
A more recent result by Bender et al. (bender2020contention, ) shows that w.h.p. constant throughput can be achieved by an adaptive algorithm with no-CD. In contrast, to the adaptive algorithm (for unknown $n$) of De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), the result by Bender et al. (bender2020contention, ) allows packets to terminate immediately upon success.
@@
In a multi-channel setting, with and without CD, Christodoulou et al. (christodoulou2018strategic, ) consider a game theoretic setting where participants may deviate from any prescribed protocol. The authors give protocols from which no participant will unilaterally deviate and which provide probabilistic latency guarantees.
Finally, recent work by Bender et al. (bender:coded, ) investigates a problem variant where at most an integer $\kappa\geq 1$ simultaneous successful transmissions are possible. This model is reminiscent of (hradovich2020contention, ) (in Section 3.1), and it captures the newer radio technology that can decode a limited number of concurrent transmissions. Here, the notion of CD alters slightly, where participants can distinguish between an empty slot and a slot with multiple transmissions; however, participants do not learn whether the number of transmissions in any slot exceeds or is at most $\kappa$ (unless they are one of the senders). Here, the authors show that a throughput of $1-o(1)$ is possible.

### v4

@@ 3.1. AQT and Contention Resolution
Bienkowski et al. (bienkowski:distributed, ; BienkowskiJKK12, ) consider a model where $n$ is known, IDs are unique, each participant has a queue, and messages may include control bits. The adversary considered here differs from the leaky-bucket and window adversaries discussed earlier: at any time, an arbitrary number of packets can be injected into the queue of any participant. The metrics of interest are the total packets in all queues, called total load, and the maximum queue size. The authors prove that their deterministic algorithm achieves a total load that is an additive $O(n^{2})$ amount larger than the best offline algorithm, and the maximum queue size is at most an $n$-factor larger.
The remaining results assume assume a model where $n$ is known, and participants have unique IDs and access to a global clock. Hradovich et al. (hradovich2020contention, ) consider a problem variant where, in any fixed round, at most $k$ stations may access the multiple access channel; this is called a $k$-restrained channel. The authors explore deterministic algorithms for leaky bucket adversaries with $\rho\leq 1$. The first is a stable adaptive algorithm whose queue size is $O(n^{2}+b)$, which is reminiscent of the bound in (chlebus2009maximum, ). The second is a full-sensing algorithm that is stable for $\rho\leq 1-1/n$. Finally, the authors give an acknowledgment-based algorithm with throughput $\Theta(k/n(\log^{2}n))$, which is complemented by a lower bound that no acknowledgement-based algorithm can achieve a throughput better than $\min\{k/n,1/(3\log n)\}$. Simulations indicate that the proposed algorithms perform favorably, with respect to queue size and the number of channel accesses, relative to well-known backoff algorithms.
@@
Summary.
There are several significant messages conveyed by the literature surveyed above. Randomized backoff algorithms continue to disappoint in that they are stable only for sub-constant injection rates. Interestingly, in terms of latency, deterministic algorithms can achieve performance that rivals and even surpasses traditional randomized backoff algorithms. However, most of these results make assumptions about unique IDs and/or a priori knowledge of $n$ (an exception is the algorithm by Anantharamu and Clebus (anantharamu2015broadcasting, ), which assumes a $1$-activating adversary). This leads to a more general message: while stability and fairness seem like natural qualities that we desire in algorithm design, it is often not possible to achieve both simultaneously when there is even a little (i.e., constant) burstiness. Therefore, all of this paints a fairly dismal picture of what is possible under window and leaky bucket adversaries, given the many impossibility results that apply.
What avenues are there for positive results? As we will discuss later in Section 3.3, new randomized approaches offer better performance. However, fundamentally, perhaps the AQT model is overly pessimistic; this has been investigated in more general settings (e.g., (BKS14, ).) Here, focusing on whether queue sizes can become unbounded under a very specific injection pattern, and under an infinite stream of packets, may lead to results that are artificially bleak. In the next section, we survey work that is concerned with a finite stream of packets that arrive under an arbitrary schedule.
@@ 3.2. Deterministic Approaches
In this section, we depart from parameterized adversaries and notions of stability, while still focusing on deterministic contention-resolution algorithms. Much of the contention-resolution work in this setting is referred to as $k$-selection, which is described as follows. There are $n$ participants of which $k$, for $1\leq k\leq n$, activate over time, each with a single packet to transmit. The activation times are selected by an adversary. As before, most of the literature operates in a model where each participant has a unique ID. Furthermore, $n$ is typically assumed to be known; throughout this section, we assume this to be true unless explicitly stated otherwise. The goal is to minimize the latency (recall from Section 1.1.3) required for all $k$ participants to successfully send their respective packets.
@@
Many results put a spin on John von Neumann’s quote by generating deterministic results from randomized approaches. Specifically, a common technique in this area is to design a randomized algorithm, and then use the probabilistic method (alon2016probabilistic, ) to argue that a deterministic analog exists. Here, we focus on the result by De Marco and Kowalski (doi:10.1137/140982763, ); however, related approaches are used in several other papers to give existence proofs for efficient transmission schedules (DBLP:conf/icdcs/MarcoKS19, ; de2013contention, ; DEMARCO20171, ; Chlebus:2016:SWM:2882263.2882514, ; ChlebusK04, ; ChlebusGKR05, ). De Marco and Kowalski (doi:10.1137/140982763, ) consider the dynamic $k$-selection problem where there is a global clock, $k$ is unknown, and no-CD. Their main result is a non-adaptive, deterministic algorithm for contention resolution with latency $O(k\log(n)\log\log(n))$.
@@
Finally, note that by the sum of a geometric series, the number of slots in $\mathcal{M}$ over which any participant executes prior to reaching row $\lg(k)+1$ is $O(k\log n\log\log n)$. Adding to this the $O(k\log(n\log n))$ slots for the Komlós and Greenberg result yields a total latency of $O(k\log n\log\log n)$.
@@ 3.2.1. Deterministic $k$-Selection
We start by reiterating that Greenberg and Winograd (greenberg1985lower, ) give a lower bound for deterministic algorithms of $\Omega(k\log n/\log k)$ latency in the static case with CD. This bound, of course, applies to the dynamic case. In particular, as discussed above, De Marco and Kowalski (doi:10.1137/140982763, ) demonstrate a non-adaptive, deterministic algorithm that achieves $O(k\log(n)\log\log(n))$ latency, which not far off of the lower bound by Greenberg and Winograd (greenberg1985lower, ).
Subsequently, De Marco et al. (DBLP:conf/icdcs/MarcoKS19, ) examine the power of acknowledgement-based versus non-acknowledgement-based deterministic non-adaptive approaches when there is no global clock and under the no-CD setting. For unknown $k$, they give a non-acknowledgement-based algorithm with latency $O(k^{2}\log n)$, and show that this can be improved to $O(k^{2}\log n/\log k)$ with an acknowledgment-based algorithm. These results are (nearly) tight; the authors also provide lower bounds on latency of $\Omega(k^{2}/\log k)$ for both cases. This illustrates a startling contrast with the case for known $k$, where the authors give an acknowledgement-based algorithm with $O(k\log k\log n)$ latency, which is known to be nearly tight (ClementiMS03, ).
@@ 3.2.2. Deterministic Wakeup
Unlike the static version, dynamic wakeup differs significantly from LE. Here, a subset of participants become awake according to some arbitrary schedule over time, while the remainder are asleep. For consistency in discussing the dynamic setting, latency (rather than makespan) is our terminology used to report results.
@@
Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider a multi-channel scenario with no-CD and no global clock. Here, all stations may access all $b\geq 1$ channels at once, and with the same parameter $k\geq 1$ being an upper bound on the number of stations that become active over time. The authors give a non-adaptive, deterministic algorithm for the cases where $k$ is unknown that has latency $O(k\log^{1/b}k\log n)$. When $b=\Omega(\log\log n)$, a faster deterministic algorithm is possible, with $O((k/b)\log n\log(b\log n))$ latency, and a lower bound of $\Omega((k/b)\log(n/k))$ is shown for deterministic algorithms. The authors also consider probabilistic jamming (a topic discussed further in Section 4), where each channel is jammed independently and uniformly with known probability $p$ per round, where $10<p<$ 1. Here, Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) give two deterministic algorithms that succeed with probability $1-O(1/\texttt{poly}(n))$, the first with latency $O(k\log(n)(\log^{1/b}k)/\log(1/p))$, and the second with latency $O((k/b)\log(n)\log(b\log(n))/\log(1/p))$ when $b=\Omega(\log(b\log n))$.
@@
Finally, under the SINR channel model, Jurdziński and Stachowiak (jurdzinski2015cost, ) show that any deterministic algorithm with access to a global clock must incur $\Omega(\log^{2}n)$ latency. This complements an earlier result by Jurdziński et al. (jurdzinski2013distributed, ), who give a deterministic algorithm with $O(\log^{2}n)$ latency in the single-hop setting.
Summary.
Much of the theoretical territory has been mapped out with some nearly-matching upper and lower asymptotic bounds on latency, and the results are less bleak than those surveyed under the AQT model. Notably, with no-CD and assuming a global clock, we see this for both $k$-selection (De Marco and Kowalski (doi:10.1137/140982763, )) and for wakeup (De Marco and Kowalski (DEMARCO20171, ; de2013contention, )), where the latter result depends on some limited system information.
Despite providing a rosier picture, deterministic approaches for contention resolution fall short of being able to offer linear latency in the dynamic case; this is unsurprising given the analogous situation in the static case (Section 2.2). Similarly, deterministic solutions to the wakeup problem yield high latency in cases without a global clock, or where key parameters are unknown. A natural place to look for improvement—perhaps even asymptotic optimality—is the domain of randomized algorithms, which we survey next.
@@ 3.3. Randomized Approaches for Contention Resolution
When a packet $u$ arrives, it listens to the channel to determine whether it is already in use by other packets. Any packets that have arrived ahead of $u$ may still be attempting to succeed and, if so, they should also be emitting a “busy signal”, which we describe momentarily. If $u$ detects this busy signal, then it listens until this signal ceases, at which point it can start attempting to transmit. In this way, there are often two groups of packets: (1) active packets that are trying to transmit, while maintaining a busy signal, and (2) idle packets that are waiting for their turn on the channel. Given this isolation, active packets can execute a contention-resolution algorithm for the static case in order to succeed; afterwards, any idle packets can become active and repeat this process.
@@
The busy signal can take different forms. For example, given CD and a global clock, active packets can emit a busy signal in each even-numbered slot; see Figure 6. Any idle packet that hears this signal or a collision infers that active packets are present and keeps waiting. However, having all packets send in every second slot may not be desirable, since sending consumes energy (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ). To reduce this cost, packets may emit the busy signal probabilistically, which can greatly reduce the sending cost; however, there is now a probability of error in the situation where some packets may still be active, but fail to emit a busy signal. When this happens, any idle packets will become active prematurely and interfere with the already-active packets.
@@
This interference is mitigated in different ways in the literature. Bender et al. (BenderKPY18, ; bender:contention, ) assume CD and $O(\log n)$ control bits (recall Section 1.1.4) in each transmission on the shared channel. The busy signal occurs probabilistically every $O(1)$ slots, with each active packet sending an expected $O(1)$ times to maintain the signal. Interleaved with the busy signal are slots used to estimate the number of active packets, $m\leq n$, which allows them to execute TSTB (recall Section 2.1) starting with a window of size $\Theta(m)$. Additionally, a leader is elected from the set of active packets to broadcast how many slots remain in the execution of TSTB; this serves as a busy signal during the execution of TSTB. The leader is frequently changed so that each packet is responsible for $O(1)$ transmissions in expectation. Newly-arrived idle packets need only listen for a handful of slots to receive this information, which allows them to (i) avoid becoming active prematurely and (ii) avoid any further listening until the current execution of TSTB is finished. While the above process can fail—for example, the estimate is poor, the busy signal fails, or active packets do not succeed in the execution of TSTB—this occurs with probability at most polynomially small in the number of active packets, which allows for $O(n)$ expected latency. Additionally, the expected number of channel accesses performed per packet is $O(\log\log^{*}n)$. This low number of accesses can be important to modern networks where devices are battery-powered, given that listening and sending on the channel consume significant energy.
@@ 3.3.1. Randomized $k$-Selection
De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ) explore randomized solutions with no-CD. When $n$ is known (or a known to within a constant multiplicative factor), they provide a non-adaptive algorithm with $O(n)$ latency, which is asymptotically optimal. For the case where $n$ is unknown, a lower bound of $\Omega(n\log n/(\log\log n)^{2})$ latency is shown to hold for any non-adaptive, randomized algorithm; thus, there is an asymptotic separation in latency between adaptive and non-adaptive randomized algorithms. Additionally, for $n$ unknown, the authors provide a strongly non-adaptive, non-acknowledgment-based algorithm with $O(n\log^{2}n)$ latency; this latency can be improved to $O(n\log^{2}n/(\log\log n))$ with an acknowledgment-based algorithm, which comes close to matching their lower bound. Lastly, when $n$ is unknown, the authors provide an adaptive algorithm that achieves $O(n)$ latency. Informally, this algorithm employs LE to break the dynamic problem into a sequence of static contention-resolution problems, as discussed earlier in this section.
A more recent result by Bender et al. (bender2020contention, ) shows that w.h.p. constant throughput can be achieved by an adaptive algorithm with no-CD. In contrast to the adaptive algorithm (for unknown $n$) of De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), the result by Bender et al. (bender2020contention, ) allows packets to terminate immediately upon success.
@@
In a multi-channel setting, with and without CD, Christodoulou et al. (christodoulou2018strategic, ) consider a game theoretic setting where participants may deviate from any prescribed protocol. The authors give protocols from which no participant will unilaterally deviate and that provide probabilistic latency guarantees.
Finally, recent work by Bender et al. (bender:coded, ) investigates a problem variant where at most an integer $\kappa\geq 1$ simultaneous successful transmissions are possible. This model is reminiscent of (hradovich2020contention, ) (in Section 3.1), and it captures newer radio technology that can decode a limited number of concurrent transmissions. Here, the notion of CD alters slightly: participants can distinguish between an empty slot and a slot with multiple transmissions, but they do not learn whether the number of transmissions in any slot exceeds or is at most $\kappa$ (unless they are one of the senders). Here, the authors show that a throughput of $1-o(1)$ is possible.